OpenStudy (anonymous):

How do you find the endpoint of R(6,-2), M(5,3)?

OpenStudy (anonymous):

from 6 to 5 is left one unit, left one more goes to 4 from -2 is 3 is up 5 units, up five more gives 8 answer is (4,8)

OpenStudy (anonymous):

x1+x2 y1+y2 _______ , _______ = Midpoint formula. 2 2 To find the EndPoint 6+x2 -2+y2 _______ = 5 ________ = 3 2 2 5*2 = 10 3*2 = 6 6+x=10 -2+y = 6 x = 10-6 y = 6 + 2 x = 4 y = 8 Therefore Endpoint = (4,8)